# Deterministic Legal Retrieval: An Action API for Querying the SAT-Graph RAG

### v1

@@ 4.2 Use Case 2: Causal Lineage Tracing for Legal Audit
This query has two distinct requirements that are impossible for standard systems: determining historical provenance ("which law introduced…") and confirming current validity with precision. A standard RAG system lacks a causal model to trace origins and cannot guarantee that a retrieved text is the most current, valid version.
@@
The agent recognizes two parallelizable sub-tasks and starts by grounding the reference to get a stable identifier.
// Initial, common step for both tasks
candidates = resolveItemReference(
reference_text="Article 227, paragraph 4 of the Constitution"
target_item_id = candidates[0].id // Agent proceeds with the top candidate
The agent can now execute the two tasks in parallel.
Task A: Determine Provenance
@@
    Retrieve the Past Version: The agent gets the specific Version that was valid on a given date in 2012.
    // Timestamp is in ISO 8601 UTC format for consistency
    version_2012 = getValidVersion(
    item_id=target_item_id, timestamp="2012-01-01T12:00:00Z"
@@
    Trace Causality to the Source: It then traces this specific version back to the action that created it.
    causality = traceCausality(version_id=version_2012.id)
    creating_action = causality.creating_action
@@
    Hydrate the Source for Readability (Agent Logic): The creating_action object contains the source_version_id. To provide a human-readable answer, the agent "hydrates" this ID (via getVersion and then getItem) and traverses its ancestors (via getAncestors) to find the root Work (the amending law) and get its label.
Task B: Determine Current State
  1. 1.
    Retrieve the Current Version: Using the same item ID, the agent fetches the version valid as of "now".
    // "NOW" is represented by the current UTC timestamp
    version_current = getValidVersion(
    item_id=target_item_id, timestamp="2025-10-06T10:30:00Z"
  2. 2.
    Compare Historical and Current Versions: To check if the text has changed, the agent compares the version from 2012 with the current one.
    diff_report = compareVersions(
    version_id_a=version_2012.id,
    version_id_b=version_current.id
    // If the diff_report is empty, the text is unchanged.
@@
The agent’s planner combines the results. From Task A, it has the identity of the amending law. From Task B, it knows whether the text has been altered. This structured data is passed to an LLM, which can now generate a highly accurate and useful answer: "The version of Article 227, paragraph 4, valid in 2012, was introduced by **[Name of Amending Law from Task A]**, effective on [date from creating_action]. A comparison with the current version shows that the text **has not changed**. It reads: ’…’". This demonstrates how an agent can compose API primitives to build a complete and auditable answer to a complex, multi-part question.
4.3 Use Case 3: Hierarchical Impact Summarization
User Query: "Provide a summary of all legislative changes, specifically creations and revocations, within the ’National Tax System’ chapter of the Constitution between January 1, 2019, and December 31, 2022."
Challenge
This query is impossible for systems that are not hierarchy-aware. A flat-text system would perform a keyword search for "tax," retrieving irrelevant articles and missing the implicit scope of the "National Tax System" chapter. Furthermore, it cannot differentiate between different types of legislative changes.
Agent Execution Plan
The agent leverages the API’s powerful server-side aggregation capabilities to construct a highly specific and efficient query plan.
  1. 1.
    Identify the Scope: The agent first resolves the reference to the chapter to get its canonical id.
    candidates = resolveItemReference(
    reference_text="National Tax System chapter of the Constitution"
    scope_id = candidates[0].id
  2. 2.
    Request a Filtered Impact Summary: The agent makes a single call to the aggregate analysis action. It constructs a request object and sends it as the body of a POST request to the summarizeImpact endpoint.
    // The agent constructs a request body for a POST call
    request_body = {
    "item_ids": [scope_id],
    "time_interval": [
    "2019-01-01T00:00:00Z",
    "2022-12-31T23:59:59Z"
    "action_types": ["Creation", "Revocation"]
    report = summarizeImpact(body=request_body)
  3. 3.
    Hydrate Details for Reporting (Agent Logic): The returned ImpactReport contains lightweight lists of identifiers. To build a rich, human-readable narrative, the agent "hydrates" these IDs into full objects using efficient batch calls.
    // The response report contains direct lists of IDs
    action_ids = report.actions
    affected_item_ids = report.affected_items
    // Get full data for the legislative events
    full_actions = getBatchActions(ids=action_ids)
    // Get full data for the affected constitutional articles
    full_items = getBatchItems(ids=affected_item_ids)
    // Additional hydration can be done to find the names of the amending laws
Synthesized Outcome
The agent now possesses a complete, structured dataset of all relevant changes. This verifiable data is passed to an LLM for synthesis. The resulting summary is not a guess, but a factual report grounded in the graph: "Between 2019 and 2022, the ’National Tax System’ chapter saw the following changes matching your criteria: Article 156-A was created by Constitutional Amendment No. 110/2019, and **Article 149-A** was revoked by Constitutional Amendment No. 109/2021." This demonstrates a complete, efficient, and auditable workflow, from high-level aggregation to detailed, human-readable reporting.
4.4 Use Case 4: Distinguishing Structural and Normative Predecessors
User Query: "What was the previous rule that governed this matter?"
Challenge
This seemingly simple question hides a deep legal ambiguity. "Previous rule" can have two distinct meanings: (1) the structural predecessor (the immediately preceding version of the same article) or (2) the normative predecessor (an entirely different article, possibly in another law, that was revoked and replaced). Standard systems can handle neither, as they lack both versioning and a semantic understanding of legislative succession.
Agent Execution Plan
Our API empowers an agent to disambiguate and solve for both interpretations. After establishing the context (getting the ‘target_item_id‘ for "this matter"), it can pursue both paths.
For the Structural Predecessor (Deterministic Causal Path):
The agent’s plan is to find the direct causal antecedent of the current version by tracing the event that created it.
  1. 1.
    It retrieves the current version of the item:
    current_version = getValidVersion(item_id=target_item_id,
    timestamp="2025-10-06T10:30:00Z", policy="SnapshotLast")
  2. 2.
    It then traces this version’s immediate cause by calling traceCausality:
    causality = traceCausality(version_id=current_version.id)
  3. 3.
    The returned CausalityTrace object contains the Action that created the current version. This Action also holds the ID of the version that it terminated. This is the structural predecessor:
    predecessor_version_id = causality.creating_action.
    terminates_version_id
This plan is fully deterministic and provides a precise, auditable answer by directly traversing the causal graph primitives.
For the Normative Predecessor (Heuristic Path):
The agent executes a more complex, discovery-oriented plan to find a semantically similar but structurally distinct predecessor.
  1. 1.
    It first gathers information about the current version: its text, and the start of its validity.
    current_version = getValidVersion(
    item_id=target_item_id, timestamp="2025-10-06T10:30:00Z", policy="SnapshotLast"
    change_date = current_version.validity_interval[0] // Start of the interval
    current_text = getTextForVersion(version_id=current_version.id, language="pt-BR").content
  2. 2.
    It then uses this information to search for semantically similar texts that were valid immediately before the change occurred.
    // Agent constructs a request body for a POST call
    // It calculates a timestamp immediately prior to the change_date
    previous_timestamp = calculate_prior_time(change_date) // e.g., change_date - 1 second
    request_body = {
    "semantic_query": current_text,
    "timestamp": previous_timestamp
    candidates = searchTextUnits(body=request_body)
  3. 3.
    The agent can then analyze the top candidates, filtering out the structural predecessor (which might also appear) and presenting the most likely normative predecessor(s) to the user.
This second plan, while heuristic, is powerfully grounded in the verifiable temporal snapshots provided by the API, making its discovery process far more reliable than a blind search.
Synthesized Outcome
The agent can now provide a sophisticated, multi-faceted answer that addresses the user’s ambiguous query head-on: "The query ’previous rule’ can have two meanings. The direct previous version of **Article 10** (its structural predecessor) had the following text: ’…’. However, this new version was created by Law XYZ, which also revoked **Article 150 of a different law** that previously governed this matter (its likely normative predecessor). Would you like to see the text of Article 150?". This demonstrates the framework’s ability to support the nuanced reasoning required for deep legal analysis.
@@ 5 Conclusion
In this paper, we addressed a critical challenge in the application of AI to the legal domain: the need for a trustworthy and auditable method for interacting with complex, versioned knowledge. While our previous work, the SAT-Graph RAG, established a robust model for representing the structural and temporal reality of legal norms, it left open the question of how to query this representation without sacrificing its inherent guarantees. We have demonstrated that the solution is not a more complex, monolithic query engine, but rather a more disciplined and transparent interaction layer.
Our primary contribution is the formal specification of the SAT-Graph API. The core of this API is a library of canonical actions—low-level, composable, and deterministic primitives that enforce auditability by design. By providing this API as the primary, formally-specified interface to the graph, we create a powerful abstraction layer. This ensures that agentic frameworks can operate on the knowledge base safely and predictably, without needing to understand the complexities of the underlying data model, making the retrieval process as rigorous as the knowledge base itself. We have shown how this architecture enables a powerful synergy: advanced agentic frameworks can provide the flexibility to interpret complex natural language queries, while our API provides the unyielding determinism required to execute the resulting retrieval plans safely. The agent’s plan becomes a complete, human-readable audit trail, transforming a potentially opaque reasoning process into a series of explicit, verifiable steps. This shift from probabilistic retrieval to deterministic, planned execution is a direct architectural response to the challenge of Explainable AI (XAI), making it possible to build trustworthy AI systems for regulated domains.
@@ 5.1 Implications
The implications of this two-layer architecture are significant, marking a paradigm shift away from end-to-end, black-box systems and toward a more modular and trustworthy approach.
@@
The most crucial implication is the decoupling of the knowledge representation from the reasoning logic. The SAT-Graph serves as the stable, verifiable "source of truth," which can be curated and maintained independently. The reasoning layer—the agentic frameworks that plan and execute queries—can evolve and improve at a rapid pace, incorporating newer and more powerful LLMs.
@@
Furthermore, our deliberate use of a generic Item data model—representing not just a document (‘Work‘) and its components (‘Work Component‘), but also abstract Concepts and Terms—means that the SAT-Graph and its accompanying API are not limited to the legal domain. This architectural pattern is directly applicable to other fields that rely on versioned, structured, and evolving documents, such as technical standards (ISO, IETF), regulatory compliance documents (FDA, financial regulations), and complex engineering manuals. In essence, any domain where what was is as important as what is can benefit from this architectural pattern.
@@
Finally, it is critical to distinguish between the responsibilities of the SAT-Graph API and those of the higher-level agentic framework that consumes it. Our API is intentionally designed as a set of verifiable primitives for retrieving structural, temporal, and causal facts, answering "what is?" and "what was?". Complex legal reasoning tasks, such as conflict detection or transitive dependency analysis, are the responsibility of the reasoning agent. This deliberate separation of concerns is a core strength of our architecture: the API provides the stable and auditable foundation of truth, while the agent provides the flexible, evolving layer of strategic reasoning.
@@ 5.2 Limitations
While this work provides a formal specification and demonstrates the conceptual power of the SAT-Graph API through use cases, several limitations should be acknowledged:
@@
: Building a production-grade implementation requires significant engineering effort, particularly for the server-side aggregation and temporal indexing capabilities.
Ontology Dependency
: The API’s effectiveness depends on the quality and completeness of the underlying SAT-Graph. Incomplete or incorrectly modeled events will propagate to retrieval results.
Agent Planning Capability
: The architecture assumes the availability of sophisticated planning agents. Current LLMs vary widely in their ability to decompose complex queries into optimal action sequences.
Evaluation Gap
: This paper focuses on specification and design. Comprehensive empirical evaluation against legal AI benchmarks remains future work.
@@ 5.3 Future Directions
This work lays a foundational layer, and in doing so, opens several avenues for future research.
First, the API specified in this paper focuses on the "verifiable" graph of hierarchical, temporal, and causal links. As highlighted by use cases like impact analysis, a critical next step is the formalization of a parallel "semantic overlay" graph. The conceptual groundwork for this involves introducing a first-class Relation entity, capable of modeling a rich vocabulary of typed, directed, and potentially time-bound links between Items (e.g., ‘cites‘, ‘succeeds‘, ‘applies_to‘). This would be exposed through a new set of ‘Graph Traversal Actions‘ in the API, enabling the tracing of transitive impacts and other complex relational queries. The practical design and formal OpenAPI specification for the Relation entity and its associated actions are being actively developed as part of an open-source project, which serves as the reference implementation for the concepts discussed in this paper. [^2] The SAT-Graph API specification project is publicly available at: https://github.com/hmartim/sat-graph-api
Second, the sophistication of the planner agent is a key area for improvement. Future work could explore fine-tuning LLMs specifically for the task of translating complex questions into optimal execution DAGs using our API. This could involve creating a synthetic dataset of questions and their corresponding gold-standard action plans for various structured document domains.
Finally, this framework can be extended to handle cross-document analysis across an entire corpus. This would involve connecting multiple SAT-Graphs and introducing actions that can trace references and dependencies not just within a single document, but between different standards, regulations, and their historical precedents, moving closer to a comprehensive model of a complete knowledge ecosystem.

### v2

@@ 4.3 Use Case 3: Thematic and Hierarchical Impact Analysis
This broad, analytical query is completely intractable for systems that are not simultaneously aware of hierarchy, themes, and time. A standard RAG system would likely perform a simple keyword search for "Digital Security," retrieving a mix of relevant and irrelevant provisions while completely missing the implicit scope of a thematic category. It cannot understand that if a high-level provisions is tagged with a theme, all its sub-components are also implicitly part of that scope. The system lacks the multi-faceted awareness to even begin to scope the query correctly, let alone analyze its evolution over time.
@@
The agent executes a sophisticated, multi-phase plan. It first deterministically resolves the complex, multi-layered scope of the query. Only after this scope is fully defined does it proceed to the historical analysis phase. The plan is designed for maximum efficiency, using parallel execution and optimized, scoped queries.
Phase 1: Deterministic Scope Resolution
@@
    Ground Thematic Scope: The agent first calls resolveThemeReference to find the canonical ID for the "Digital Security" theme. It then expands this scope by calling a navigation primitive to get all descendant themes.
    resolveThemeReference(reference_text="Digital Security")
    // Followed by getThemeHierarchy(...)
@@
    Find Anchor Items: The agent performs a single, highly efficient searchItems call that retrieves all items that are linked to the ’Digital Security’ thematic tree. This efficiently finds the high-level "anchor" items.
@@
    Hierarchical Expansion: For each anchor item found, the agent makes a parallel call to getItem Hierarchy to gather the IDs of all its descendants. This correctly propagates the thematic relevance down the hierarchy (e.g., from a tagged article to all its untagged paragraphs). The union of these sets forms the final, complete list of relevant provision IDs.
Phase 2: Parallel Historical Analysis
  1. 4.
    Fetch Historical Events: Now, with a precise list of hundreds of item IDs, the agent avoids making individual history calls. It strategically groups the IDs by their root document (e.g., all from the 1988 Constitution). For each group, it makes a single, powerful API call to retrieve all Action objects that affected those specific items since the year 2000. These calls are executed in parallel for maximum performance.
  2. 5.
    Aggregate and Synthesize: The agent collects the results from all parallel calls, now possessing a complete, verifiable dataset of every relevant change. This structured data is aggregated to produce statistics (e.g., 5 amendments, 2 revocations) and then passed to an LLM.
@@
The LLM receives a rich, structured context containing every single piece of evidence. It can then generate a high-level summary grounded in facts, not probabilistic guesses. The response can be both quantitative and narrative: "Since 2000, there have been X legislative actions affecting Y provisions related to ’Digital Security’. The most significant change was Constitutional Amendment No. 115/2022, which elevated data protection to a fundamental right by amending Article 5…" Each claim in this summary can be traced back to a specific Action object retrieved by the agent, making the entire analytical process fully auditable and trustworthy.
@@ 5 Conclusion
In this paper, we addressed a critical challenge preventing the safe deployment of autonomous AI in the legal domain: the lack of a trustworthy and auditable method for interacting with complex, versioned knowledge. While knowledge graphs like our SAT-Graph can represent the structural and temporal reality of legal norms, the question of how an agent can query this representation without sacrificing its guarantees remained open. We have demonstrated that the solution is not a more complex, monolithic query engine, but rather a more disciplined and transparent interaction layer: a formal Canonical Primitive API.
Our primary contribution is a new architectural pattern for building deterministic legal agents, specified through this API. The core of the framework is a library of canonical primitives—low-level, composable primitives that enforce auditability by design. This approach creates a powerful synergy: advanced agentic frameworks provide the flexibility to interpret natural language, while our API provides the unyielding determinism required to execute the resulting reasoning plans safely. The agent’s plan itself becomes a complete, human-readable audit trail, transforming its internal reasoning process from an opaque black box into a series of explicit, verifiable steps.
We have shown that this shift—from passive, probabilistic retrieval to active, agent-driven inquiry built on deterministic primitives—is a direct architectural response to the challenge of Explainable AI (XAI). It provides a concrete blueprint for building the next generation of trustworthy AI systems, not just for law, but for any regulated domain where precision, auditability, and historical correctness are paramount.
@@ 5.1 Implications
The implications of enforcing a formal Primitive API between an agent and a temporal knowledge graph are significant, marking a paradigm shift away from end-to-end, black-box systems and toward a more modular and trustworthy approach.
@@
The most essential implication is the clean decoupling of the knowledge representation (the graph) from the agentic reasoning layer. The SAT-Graph serves as the stable, verifiable "source of truth," which can be curated and maintained independently. The reasoning layer—the agentic frameworks that plan and execute queries—can evolve at a rapid pace, incorporating newer and more powerful LLMs without compromising the foundational integrity of the knowledge base.
@@
This Primitive API architecture is not a solution only for law; it is a generalizable pattern for building trustworthy agents in any regulated, structured domain. Our deliberate use of a generic Item data model means that this pattern is directly applicable to other fields that rely on versioned and evolving documents, such as technical standards (ISO, IETF), regulatory compliance (FDA, financial regulations), and complex engineering manuals. In essence, any domain where what was is as important as what is can benefit from this approach to enabling auditable agentic systems.
@@
Finally, it is critical to distinguish between the responsibilities of the Primitive API and those of the higher-level agentic framework that consumes it. Our API is intentionally designed as a set of verifiable primitives for retrieving structural, temporal, and causal facts. It answers "what is?" and "what was?". Complex reasoning tasks, such as conflict detection or transitive dependency analysis, remain the responsibility of the reasoning agent. This deliberate separation of concerns is a core strength of our architecture: the API guarantees the integrity of the evidence, while the agent is responsible for constructing the argument.
@@ 5.2 Limitations
While this work proposes a complete architectural pattern and formal API specification, several limitations and areas for future work should be acknowledged:
@@
Building a production-grade implementation of the server-side logic is a significant engineering effort. The performance of the API, particularly for temporal and causal queries over a large graph, depends on sophisticated indexing strategies that are non-trivial to implement and optimize.
Knowledge Base Dependency
The determinism of the agent’s reasoning is fundamentally bound to the quality and completeness of the underlying knowledge graph. The principle of "garbage in, garbage out" applies: incomplete or incorrectly modeled legislative events in the SAT-Graph will lead to factually correct but legally incomplete answers from the agent. The API provides transactional integrity but cannot create knowledge that is not there.
Agent Planner Sophistication
Our architecture fundamentally shifts the primary challenge from retrieval to agentic planning. The effectiveness of the entire system now relies on the agent’s ability to translate a user’s natural language query into an optimal and correct sequence of primitives. While the capabilities of LLMs as planners are advancing rapidly, generating complex, multi-step, and efficient DAGs of primitives remains a frontier research problem.
Empirical Evaluation Gap
This paper focuses on the formal specification and conceptual power of the architecture. A comprehensive empirical evaluation against emerging legal AI benchmarks that test temporal and causal reasoning is a critical next step. Such an evaluation would be necessary to quantify the performance and accuracy gains of this agentic, primitive-based approach over traditional RAG models.
@@ 5.3 Future Directions
By establishing a formal, primitive-based interface for agentic reasoning, this work lays a foundational layer for numerous avenues of future research.
First, we plan to expand the API’s "grammar" to support deeper legal reasoning. The current specification focuses on the verifiable graph of structural, temporal, and causal links. A critical next step is the formalization of a parallel "semantic overlay" graph. This involves introducing a first-class Relation entity to model a rich vocabulary of typed links between Items (e.g., ‘cites‘, ‘succeeds‘, ‘applies_to‘). A new set of ‘Graph Traversal Primitives‘ would then empower an agent to trace transitive impacts and other complex relational queries, moving from "what happened?" to "how are these concepts legally related?". The design for these extensions is being actively developed in our open-source project. [^4] The SAT-Graph API specification project is publicly available at: https://github.com/hmartim/sat-graph-api
Second, as noted in our limitations, the sophistication of the planner agent is now a key determinant of the system’s overall capability. The next challenge is to improve the agent’s fluency in translating complex questions into optimal execution DAGs using our API. Future work will explore fine-tuning LLMs specifically for this "text-to-plan" task. This involves creating a synthetic dataset of natural language questions and their corresponding gold-standard primitive plans, which can be used to teach agents how to reason effectively over structured, temporal domains.